# Bullet: Boosting GPU Utilization for LLM Serving via Dynamic Spatial-Temporal Orchestration

### v1

@@ Abstract.
Modern LLM serving systems confront inefficient GPU utilization due to the fundamental mismatch between compute-intensive prefill and memory-bound decode phases. While current practices attempt to address this by organizing these phases into hybrid batches, such solutions create an inefficient tradeoff that sacrifices either throughput or latency, leaving substantial GPU resources underutilized. We identify two key root causes: 1) the prefill phase suffers from suboptimal compute utilization due to wave quantization and attention bottlenecks. 2) hybrid batches disproportionately prioritize latency over throughput, resulting in wasted compute and memory bandwidth. To mitigate the issues, we present Bullet, a novel spatial-temporal orchestration system that eliminates these inefficiencies through precise phase coordination. Bullet enables concurrent execution of prefill and decode phases, while dynamically provisioning GPU resources using real-time performance modeling. By integrating SLO-aware scheduling and adaptive resource allocation, Bullet maximizes utilization without compromising latency targets. Experimental evaluations on real-world workloads demonstrate that Bullet delivers 1.26 average throughput gains (up to 1.55) over state-of-the-arts, while consistently meeting latency constraints.
@@ 1. Introduction
GPUs are the predominant computing platform for large language model (LLM) services (OpenAI, 2023; DeepSeek-AI et al., 2024; Yang et al., 2024; Dubey et al., 2024), empowering diverse applications with varying computational and latency requirements (Hong et al., 2024a; lla, 2022; Hong et al., 2024b). As these applications continue to advance, maximizing GPU utilization has become crucial for elevating serving quality (Zhu et al., 2024; Lin et al., 2024). In response, a plethora of systems have been developed to optimize key aspects of LLM serving, such as kernel performance (Kwon et al., 2023; Dao et al., 2022; Prabhu et al., 2025), scheduling strategies (Agrawal et al., 2023a; Zhong et al., 2024; Patel et al., 2024), and parallelization techniques (Wu et al., 2024; Qin et al., 2025).
@@
However, given the divergent computational characteristics in LLM inference workflow, achieving high resource utilization poses significant challenges (Zhu et al., 2024). In detail, the workflow involves a computationally intensive prefill phase that processes all inputs in parallel, succeeded by a memory-bound decode phase that generates tokens sequentially. These two contrasting phases lead to an imbalanced utilization of computational resources and memory bandwidth. The dynamism of workload patterns further widens the gap between the two phases, enforcing GPUs to switch between these complementary resource utilization states. Moreover, the distinct computational characteristics between the two phases introduces an unavoidable throughput-latency tradeoff (Agrawal et al., 2023a). For example, conventional systems (Kwon et al., 2023) (Figure 1a) prioritize the prefill, which subsequently leads to a larger decode batch size. As a result, the strategy enhances throughput but incurs decode latency penalties, as the prefill operations exclusively occupy the GPU resources. Therefore, an ideal system necessitates fine-grained control to balance throughput, latency and GPU utilization, are non-trivial to achieve.
@@
Prior attempts either isolate the two phases across different GPUs or coordinate in the same devices with spatial multiplexing. Prefill-decode disaggregation (Zhong et al., 2024; Patel et al., 2024; Gao et al., 2024; DeepSeek-AI et al., 2024; Qin et al., 2025) implements the first strategy by physically separating the phases across dedicated GPU groups. However, such systems require careful determination of GPU counts for each phase, tailored to specific workload patterns (DeepSeek-AI et al., 2024), and may lack rapid reconfigurations for fluctuating request patterns (Wu et al., 2024). Moreover, high-bandwidth interconnects are critically demanded to handle the significant state migration overhead (Qin et al., 2025; Zhong et al., 2024), greatly limiting the scenarios for deployment. In practice, chunked prefill (Agrawal et al., 2023a, b) (Figure 1b) has been pervasively employed in production systems (Kwon et al., 2023; Zheng et al., 2023; NVIDIA, 2023; Team, 2025). This technique leverages a fixed token budget to combine prefill and decode requests into hybrid batches, with longer sequences being partitioned into chunks to fit within capacity. While tunable chunk sizes enable latency control, the benefits inevitably compromise throughput (Agrawal et al., 2023a; Prabhu et al., 2025; Zhu et al., 2024). The optimization presents inherent tensions, where smaller chunks sizes may operate below GPU saturation points, and larger ones sacrifice latency enhancements.
Despite throughput-latency tradeoff in LLM serving has been extensively studied (Agrawal et al., 2023a; Prabhu et al., 2025), we identify a critical limitation that, these approaches often lead to GPU under-utilization, resulting in suboptimal point stressing the balance of latency and hardware efficiency. First, the prefill phase achieves at most 75% of peak sustainable compute utilization regardless of sequence length. For short sequence, severe wave quantization effects (Corporation, [n. d.]) leaves an average of 19% computation unit cycles idle, while for long sequences, attention operations (Kwon et al., 2023; Dao et al., 2022), which maintain lower GPU utilization than linear layers dominates execution time. Second, the throughput-latency tradeoff in chunked prefill exhibits sub-linearly scaling with chunk sizes. Modest reductions in chunk size disproportionately degrade performance while yielding diminishing latency benefits, creating an inefficient optimization landscape. The key reason is that prefill and decode sequences are tightly coupled in the hybrid batches and executed in lock-step that under-utilizes resources, with an average of 64% compute and 16% memory bandwidth. Third, spatial sharing should properly leverage the natural complementarity between compute-intensive prefill and memory-bound decode phases through concurrent execution (Figure 1c) with minimal synchronization. However, the system still demands carefully orchestrating the concurrent phases to reduce contention and potential latency violation.
To mitigate the under-utilization issue and balance throughput-latency tradeoff, we propose Bullet, an LLM serving system to saturate GPU resources through spatial-temporal orchestration with precise provision. The system proactively monitors request progress and rapidly adjusts resource provision to maintain high utilization while adhering to latency requirements. Bullet achieves such efficient execution with four key components: 1) performance estimator (section 3.2) that establishes a profile-augmented analytical model; 2) SLO-aware task scheduler (section 3.3) to dynamically balance prefill and decode phases; 3) computational resource manager (section 3.4) that offers lightning yet precise resource configuration; 4) concurrent execution engine (section 3.5) for asynchronous CPU control flow and GPU execution.
@@
We identify the inefficiencies in maintaining high GPU utilization while navigating throughput-latency tradeoffs in existing LLM serving systems, and systematically analyze the chances for boosting the utilization.
We establish accurate modeling for spatial-temporal shared phases, and propose fine-grained control to optimize resource provision aware of latency requirements.
We develop an LLM serving system to effectively integrate the proposed techniques onto readily available frameworks.
Experimental evaluations on real-world workloads show that Bullet outperforms state-of-the-arts in both latency and throughput while achieving higher GPU utilization.
@@ 2.1. LLM Computational Workflow
Recent LLMs (OpenAI, 2023; Dubey et al., 2024; Yang et al., 2024) are mainly built by stacking Transformer blocks (Vaswani et al., 2017), with each containing four core components: QKV-projection layer, self-attention computation, Output-projection layer and multi-layer perception (MLP). These operators are primarily implemented through general matrix multiplications (GEMMs), with element-wise operations interspersed between layers. The self-attention specifically operates on query, key, and value matrices generated by the QKV-projection, with the computational efficiency significantly enhanced through optimized kernel implementations like FlashAttention (Dao et al., 2022).
The LLM inference pipeline operates through two distinct computational phases. During the initial prefill phase, the system processes all input tokens in parallel to generate the first output token, with the latency measured as time-to-first-token (TTFT). This compute-intensive stage performs full attention computations across the entire sequence while building the KV cache to store intermediate key-value states. Subsequently, the decode phase generates tokens sequentially, with each iteration consuming only the most recent output token to produce the next. The average iteration latency is termed time-per-output-token (TPOT). This memory-bound phase primarily reads from the KV cache while performing minimal computation for the current token’s transformations.
@@ 2.2. GPU Utilization Principles
Sequence Length | QKV | Attn | OProj | MLP | Total |
@@
2.2.1. Execution Model and Theoretical Performance Bound.
During execution, multiple TBs can reside on the same SM to share the hardware resources. The SM executes the warps (32-thread groups) in a wavefront fashion to interleave different instructions and maximize hardware utilization. However, if the number of TBs is not evenly divisible by the number of SMs, a workload imbalance situation called wave quantization (Osama et al., 2023; Corporation, [n. d.]) occurs. Some SMs finish early and remain idle while waiting for others to complete the tail wave. Formally, given a grid size of TBs and a GPU with SMs, the kernel demands waves to complete, with only SMs active in the final wave. Equation 1 computes the idle SM cycle ratio to the total number of cycles in a kernel.
GPU kernels typically use power-of-2 grid sizes to match data dimensions, but this conflicts with non-power-of-2 SM counts in GPUs. For example, 108 for Nvidia A100 (Corporation, 2021) and 132 for H100 (Corporation, 2022). Therefore, wave quantization remains an open issue (Osama et al., 2023; Jangda et al., 2024) across diverse kernels. This inefficiency is particularly pronounced in Transformer’s self-attentions (Dao et al., 2022; Vaswani et al., 2017) and small-shaped GEMMs for short input sequences or small chunked prefill-sizes (section 2.3.1) (Agrawal et al., 2023a). Additional under-utilization cause arises from memory-bound kernels like LLM’s decode phases and element-wise operators, which idle compute resources during frequent memory accesses.
2.2.2. GPU Resource Sharing.
Naturally, compute- and memory- bound kernels are suitable to co-execute on GPU, saturating both compute and bandwidth resources (Zhang et al., 2025a; Han et al., 2022; Lin et al., 2023; Chen et al., 2022). The complementary nature of the prefill and decode phases makes them ideal for such concurrent exection. Since LLM serving systems generally necessitate adherence to service-level objectives (SLOs) of predefined latency requirements (Zhong et al., 2024; Patel et al., 2024; Agrawal et al., 2023a), predictable and controllable execution time over the two phases is demanded. However, current GPUs lack deterministic concurrent scheduling controls (Jain et al., 2019; Olmedo et al., 2020; Gilman et al., 2020; Bakita and Anderson, 2024), forcing users to carefully orchestrate kernel co-location to achieve reliable overlap (Zhang et al., 2025a; Strati et al., 2024; Zhang et al., 2025b). Although modern GPUs provide resource partitioning through Nvidia’s multi-process service (MPS) (NVIDIA, 2025), these mechanisms still require precise kernel management to ensure effective resource sharing while meeting SLO requirements.
2.2.3. Kernel-Level Analysis
@@
To characterize LLM serving performance and resource utilization, we analyze the execution time (Figure 2(a)), hardware utilization (Figure 2b,c) and theoretical waste caused by wave quantization (Table 1). The profiling methodology is detailed in section 4.1, with CPU overheads excluded. While MLP operations achieve up to 92% compute utilization, complete Transformer layers sustain only 70%-76% due to compounding inefficiencies. For shorter sequences, severe wave quantization in GEMMs creates substantial under-utilization, as evidenced by the O-projection’s measured 49% and 70% utilization, respectively. This result closely agrees with the theoretical bound of 59% and 79% ( in Table 1). Furthermore, PagedAttention’s (Kwon et al., 2023) irregular memory access patterns degrade attention kernel performance (Dao et al., 2022), which is particularly problematic as attention dominates 34% of execution time for longer sequences. Together, the effects of wave quantization and memory access inefficiencies create persistent performance gaps between theoretical peak and achieved throughput. These gaps remain substantial regardless of input sequence length, inherently constraining overall system efficiency.
@@
Takeway 1. GPUs remain underutilized even during compute-intensive prefill. While co-locating prefill and decode saturate the resources, predictable control to orchestrate the two phases is demanded.
2.3. Biased Utilization-Latency Tradeoff
@@
As shown in Figure 3a, chunked prefill (Agrawal et al., 2023a) achieves low and stable TPOT by leveraging a fixed token budget to concatenate the prefill and decode tokens into a hybrid batch, executing in a lockstep fashion (❷). Given a chunk size of , the hybrid batch is first filled with active decode requests first, and allocates the remaining tokens to the prefill sequences. Sequences exceeding this residual capacity are split into chunks, leaving residual tokens processed in subsequent iterations. Therefore, the prefill completion requires iterations. This forces times KV cache reloads as each new chunk must reprocess previous chunks’ cached states.
Due to the lock-step execution of the hybrid batch, a smaller chunk size effectively decreases TPOT at the cost of increased TTFT and degraded system throughput (Agrawal et al., 2023a), while larger chunks exhibit the opposite effect. Previous works (Agrawal et al., 2023a; Zhu et al., 2024) recognize these inherent tradeoffs and propose tuning chunk size based on workload’s prefill-to-decode (P:D) time ratio through manual tuning (Agrawal et al., 2023a; Prabhu et al., 2025; Zhu et al., 2024) or automatic searching (Agrawal et al., 2024; Cheng et al., [n. d.]; Team, 2025).
@@
Despite the throughput-latency tradeoff has been extensively studied, we highlight that such tradeoff is biased, and the resulting performance degradation remains overlooked. First, as discussed in section 2.2.3, chunked prefill typically uses suboptimal chunk sizes below GPU-saturating levels to prioritize low latency. This produces severe wave quantization effects (Corporation, [n. d.]), creating GPU bubbles (Figure 3a-❶). Second, redundant KV cache reloads required for long sequences significantly prolong attention computation time (❸), further reducing GPU utilization. Third, these factors collectively inflate TTFT, triggering a cascading congestion effect in which queued requests stall while awaiting prefill completion, degrading overall system throughput.
@@
Figure 4 systematically quantify the performance degradation of chunked prefill of a 16k-token sequence prefill even without hybrid batching. For 1k chunk size, a progressive 10% drop in compute efficiency (from 71% to 61%) across successive chunks is witnessed in Figure 4(a), which falls substantially below the 77% achievable peak. This under-utilization stems from redundant KV cache reloading in chunked attention, which also causes the final chunk’s processing time to 1.9 than that of the initial chunk. Consequently, per-chunk latency scales linearly by chunk counts and increases total prefill latency by 1.13 compared to unchunked execution. While a larger chunk size of 2k partially mitigates utilization drops from -18% to -7%, the average per-chunk latency increases by 1.86, significantly diminishing the TPOT improvements that motivated chunked prefill. This fundamental tension between maintaining high hardware utilization and minimizing TPOT creates an intractable optimization in chunked prefill.
Takeaway-2 Chunked prefill inherently falls short in achieving high utilization due to 1) small chunk size that disproportionately trades throughput for latency. 2) Redundant memory access. 3) cascading congestion on pending requests.
2.4. Existing Optimizations
Canonical chunked prefill (Agrawal et al., 2023a) sequentially computes decode and prefill attention, leading to severe bubbles. While PodAttention (Prabhu et al., 2025) mitigates this inefficiency by fusing the two attention kernels, such operator-specific optimization fails to address the inherent limitations of chunked prefill (section 2.3.2). Nanoflow (Zhu et al., 2024) optimizes the hybrid batch through finer-grained nano batches (Figure 3b). The system establishes a fixed-size processing pipeline through careful configuration of the kernel’s grid size and inter-kernel synchronization. Using CUDA streams (Corporation, 2025), the technique achieves partial overlap among compute-intensive, memory-bound, and network communication operators across distinct nano batches. However, similar to chunked prefill, the attention duration grows by successive chunks. This gap eventually eliminates the chances of overlapping due to kernel dependencies and fixed pipelines.

### v2

@@ Abstract.
Modern large language model (LLM) serving systems confront inefficient GPU utilization due to the fundamental mismatch between compute-intensive prefill phase and memory-bound decode phase. While current practices attempt to address this by organizing these phases into hybrid batches, such solutions create an inefficient tradeoff that sacrifices either throughput or latency, leaving substantial GPU resources underutilized. For this, we identify two key root causes: 1) the prefill phase suffers from suboptimal compute utilization due to wave quantization and attention bottlenecks, and 2) hybrid batching disproportionately prioritizes latency over throughput, wasting both compute resources and memory bandwidth. To mitigate the issues, we present Bullet, a novel spatial-temporal orchestration system that eliminates these inefficiencies through fine-grained phase coordination. Bullet enables concurrent execution of prefill and decode requests, while dynamically provisioning GPU resources based on real-time performance modeling. By integrating SLO-aware scheduling and adaptive resource allocation, Bullet maximizes GPU utilization without compromising latency targets. Experimental evaluations on real-world workloads demonstrate that Bullet delivers 1.26 average throughput gains (up to 1.55) over state-of-the-arts, while consistently meeting latency constraints.
@@ 1. Introduction
GPUs have become the predominant computing platform for large language model (LLM) services (OpenAI, 2023; DeepSeek-AI et al., 2024; Dubey et al., 2024), powering a wide range of applications with varying computational and latency demands (Hong et al., 2024; lla, 2022). As these applications continue to grow in scale and complexity, maximizing GPU utilization has become crucial for elevating service quality (Zhu et al., 2025; Lin et al., 2024). In response, a plethora of systems have been developed to optimize different aspects of LLM serving, such as kernel-level performance (Kwon et al., 2023; Dao et al., 2022; Prabhu et al., 2025), scheduling strategies (Agrawal et al., 2023; Zhong et al., 2024; Patel et al., 2024), and parallelization techniques (Wu et al., 2024).
@@
However, the divergent computational characteristics of LLM inference make high GPU resource utilization particularly challenging (Zhu et al., 2025). In detail, the workflow consists of a computationally intensive prefill phase that processes all inputs in parallel, succeeded by a memory-bound decode phase that generates tokens sequentially. These two contrasting phases lead to a structural imbalance in the utilization of computational resources and memory bandwidth. The dynamic nature of workload patterns further exacerbates this imbalance, forcing GPUs to alternate between these complementary resource utilization states. Moreover, the distinct characteristics of the two phases inherently introduce a throughput-latency tradeoff (Agrawal et al., 2023). For example, conventional systems (Kwon et al., 2023) (Figure 1a) prioritize the prefill, which thus increases the size of subsequent decode batches. While this improves throughput, it also penalizes latency, since prefill operations monopolize GPU resources. Therefore, an ideal system necessitates fine-grained control to balance throughput and latency without elevated GPU utilization, and is non-trivial to achieve.
@@
Prior attempts have sought to address phase imbalance either by isolating prefill and decode across different GPUs, or by coordinating them on the same device through spatial multiplexing. Prefill-decode disaggregation (Zhong et al., 2024; Patel et al., 2024; DeepSeek-AI et al., 2024) represents the first approach, physically separating the two phases across dedicated GPU groups. However, such systems require careful tuning of GPU allocations for each phase, tailored to specific workload patterns (DeepSeek-AI et al., 2024), and struggle to reconfigure quickly under fluctuating request loads (Zhong et al., 2024). Furthermore, they impose heavy demands on high-bandwidth interconnects to support frequent state migration (Qin et al., 2025), significantly restricting the deployment scenarios. In practice, chunked prefill (Agrawal et al., 2023) (Figure 1b) has been pervasively adopted in production systems (Kwon et al., 2023; Zheng et al., 2023; NVIDIA, 2023; Team, 2025). This method leverages a fixed token budget to combine prefill and decode requests into hybrid batches, with longer sequences being partitioned into chunks to fit within capacity. While tunable chunk sizes provide a degree of latency control, the approach inevitably compromises throughput (Agrawal et al., 2023; Prabhu et al., 2025). Smaller chunk sizes often underutilize GPU capacity, while larger ones negate latency improvements, reflecting the inherent trade-off in balancing efficiency and responsiveness.
Although throughput-latency tradeoff in LLM serving has been extensively studied (Agrawal et al., 2023; Zhong et al., 2024), we identify a critical limitation that, existing approaches often underutilize GPUs, resulting in suboptimal balance between latency and hardware efficiency. First, the attention exhibits relatively low compute utilization compared to linear layers (section 2.2) despite optimized implementations (Dao et al., 2022). Second, the throughput-latency tradeoff in chunked prefill exhibits sub-linear scaling with chunk sizes that prolongs execution time for successive chunks (section 2.3). Third, while spatial sharing has the potential to exploit the natural complementarity between compute-intensive prefill and memory-bound decode phases through concurrent execution (Figure 1c), effective orchestration is required. Without careful coordination, contention and latency violations remain significant risks. These observations underscore the need for dynamic coordination to fully exploit GPU performance. To address GPU under-utilization and balance the trade-off between throughput and latency, we propose Bullet, an LLM serving system that saturates GPU resources through spatial-temporal orchestration with fine-grained resource provisioning. Bullet proactively monitors request progress and dynamically adjusts resource provision to sustain high utilization while satisfying latency requirements. The system achieves such efficient execution with four key components: 1) a performance estimator (section 3.2) with an accurate analytical model of low profile and runtime overhead; 2) an SLO-aware task scheduler (section 3.3) that dynamically adjusts prefill and decode requests to balance throughput and latency; 3) a computational resource manager (section 3.4) that offers lightning yet precise resource configuration; 4) a concurrent execution engine (section 3.5) that enables asynchronous CPU control flow and GPU execution.
Bullet achieves optimal resource provisioning via an SM-scaling roofline model that captures contention between co-executed prefill and decode kernels with varying computational resource partitions. Calibrated with minimal samples, the model is capable of continuous refinement with online data. Layer-wise prefill scheduling monitors request progress and SLO compliance, reconfiguring resources through lightweight compute-unit masking (Bakita and Anderson, 2023). With microsecond-level operational overhead, Bullet maintains both precise control and high efficiency throughout execution. Bullet is available at https://github.com/zejia-lin/Bullet.
@@
We identify the inefficiencies that hinder GPU utilization in existing LLM serving systems while navigating throughput-latency tradeoffs, and systematically analyze opportunities for boosting utilization.
We establish accurate modeling for spatial-temporal shared phases, and introduce fine-grained control mechanisms for latency-aware resource provisioning.
We design and implement an LLM serving system to effectively integrate the proposed techniques into readily available frameworks.
Experimental evaluations on real-world workloads show that Bullet outperforms state-of-the-arts in both latency and throughput while achieving significantly higher GPU utilization.
@@ 2.1. LLM Computational Workflow
Recent LLMs (OpenAI, 2023; Dubey et al., 2024; Yang et al., 2025) are mainly built by stacking Transformer blocks (Vaswani et al., 2017), with each containing four core components: QKV-projection layer, self-attention computation, Output-projection layer and multi-layer perception (MLP). These components are primarily implemented through general matrix multiplications (GEMMs), with element-wise operations interspersed between layers. Among them, the self-attention specifically operates on query, key, and value matrices produced by the QKV-projection. Its computational efficiency has been significantly enhanced through optimized kernel implementations like FlashAttention (Dao et al., 2022).
The LLM inference pipeline operates through two distinct computational phases. During the initial prefill phase, the system processes all input tokens in parallel to generate the first output token, with the latency measured as time-to-first-token (TTFT). This compute-intensive stage performs full attention computations across the entire sequence while building the KV cache to store intermediate key-value states. Subsequently, the decode phase generates tokens sequentially, with each iteration consuming only the most recent output token to produce the next. The average iteration latency is termed time-per-output-token (TPOT). Unlike prefill, this memory-bound phase primarily retrieves data from the KV cache and performs relatively lightweight computations for the current token’s transformations.
@@ 2.2. GPU Utilization Principles
2.2.1. Execution Model and Theoretical Performance Bound.
During execution, multiple TBs can reside on the same SM to share its registers, shared memory, and thread slots. The TBs per SM can be obtained via hardware vendor’s runtime APIs (Corporation, 2025; hip, 2025). The SM executes the warps (32-thread groups) in successive wavefronts to interleave different instructions and maximize hardware utilization. However, if the number of TBs is not evenly divisible by the number of SMs, a workload imbalance situation called wave quantization (Corporation, [n. d.]; Lee et al., 2025) occurs. Some SMs finish early and remain idle while waiting for others to complete the tail wave. Formally, given a kernel with TBs, SMs, and TBs per SM, the kernel demands waves to complete. In the final wave, the TBs are distributed unevenly due to Most-Room Policy (Gilman et al., 2020; Bakita and Anderson, 2025), resulting in only SMs active. Therefore, the corresponding ratio of idled SM cycles can be quantified as .
GPU kernels typically use power-of-2 grid sizes to match data dimensions, but this conflicts with non-power-of-2 SM counts in GPUs. For example, 108 for Nvidia A100 (Corporation, 2021) and 132 for H100 (Corporation, 2022). Therefore, wave quantization remains an open issue (Osama et al., 2023; Jangda et al., 2024) across diverse kernels. This inefficiency is particularly pronounced in Transformer’s self-attention (Vaswani et al., 2017; Dao et al., 2022) and small-shaped GEMMs for short input sequences or small chunked prefill-sizes (section 2.3.1) (Agrawal et al., 2023). Additional under-utilization cause arises from memory-bound kernels like LLMs’ decode phases and element-wise operators, which idle compute resources during frequent memory accesses.
Seq. Len. | QKV | Attn | O | MLP | Layer’s Total |
@@
2.2.2. LLM Kernel Characteristics
@@
We quantify LLM serving efficiency by analyzing execution time (Figure 2(a)), hardware utilization (Figure 2b,c), and the theoretical waste caused by wave quantization (Table 1). While MLP operations achieve up to 92% compute utilization, complete Transformer layers sustain only 70%-76% due to compounding inefficiencies. For shorter sequences, severe wave quantization in GEMMs creates substantial underutilization, as evidenced by the O-proj’s measured 49% and 70% utilization, respectively. This result closely agrees with the theoretical bound of 59% and 79% ( in Table 1). The attention kernels, even with optimized implementations (Dao et al., 2022), exhibit lower utilization than GEMM and account for 34% of the runtime on long sequences. Together, the effects of wave quantization and attention bottleneck create persistent performance gaps between theoretical peak and achieved throughput. These gaps remain substantial regardless of sequence length, inherently constraining overall system efficiency.
2.2.3. GPU Resource Provisioning and Sharing.
Naturally, compute- and memory-bound kernels are suitable to co-execute on GPUs, saturating both compute and bandwidth resources (Han et al., 2022; Strati et al., 2024; Choi et al., 2022; Zhang et al., 2025a). The complementary nature of the prefill and decode phases makes them ideal for such concurrent execution. Since LLM serving systems generally necessitate adherence to service-level objectives (SLOs) of predefined latency requirements (Zhong et al., 2024; Patel et al., 2024; Agrawal et al., 2023), predictable and controllable execution time over the two phases is demanded. However, current GPUs lack deterministic concurrent scheduling controls (Jain et al., 2019; Olmedo et al., 2020; Bakita and Anderson, 2024), forcing users to carefully provision resources for kernels to achieve reliable overlap (Zhang et al., 2025a; Strati et al., 2024; Zhang et al., 2025b). While modern GPUs provide compute resource partitioning through Nvidia’s multi-process service (MPS) (NVIDIA, 2025), precise kernel management is still required to ensure effective resource sharing while meeting SLO requirements. Figure 3 demonstrates that prefill scales near-linearly with SM count, while decode exhibits super-linear scaling. This suggests potential throughput gains from concurrent execution with properly balanced SM allocation.
@@
Takeaway 1. GPUs remain underutilized even during compute-intensive prefill. While co-locating prefill and decode saturates the resources, precise resource provisioning to orchestrate the two phases is demanded.
2.3. Biased Throughput-Latency Tradeoff
@@
As shown in Figure 4, chunked prefill (Agrawal et al., 2023) achieves low TPOT by leveraging a fixed token budget to concatenate the prefill and decode tokens into a hybrid batch, executing in a lockstep fashion (❷). Given a chunk size of , the hybrid batch is first filled with active decode requests first, and allocates the remaining tokens to the prefill sequences. Sequences exceeding this residual capacity are split into chunks, leaving residual tokens processed in subsequent iterations. Therefore, the prefill completion requires iterations. This forces times KV cache reloads as each new chunk must reprocess previous chunks’ cached states.
Due to the lock-step execution of the hybrid batch, a smaller chunk size effectively decreases TPOT at the cost of increased TTFT and degraded system throughput (Agrawal et al., 2023), while larger chunks exhibit the opposite effect. Previous works (Agrawal et al., 2023; Zhu et al., 2025) recognize these inherent tradeoffs and propose tuning chunk size based on workload’s prefill-to-decode time ratio through manual tuning or automatic searching (Agrawal et al., 2024; Cheng et al., [n. d.]).
@@
Despite the throughput-latency tradeoff has been extensively studied, we highlight that such a tradeoff is biased, and the resulting performance degradation remains overlooked. First, as discussed in section 2.2.2, chunked prefill typically uses suboptimal chunk sizes below GPU-saturating levels to prioritize low latency. This produces severe wave quantization effects (Osama et al., 2023), creating GPU bubbles (Figure 4-❶). Second, redundant KV cache reloads required for long sequences significantly prolong attention computation time (❸), further reducing GPU utilization. Third, these factors collectively inflate TTFT, triggering a cascading congestion effect in which queued requests stall while awaiting prefill completion, degrading overall system throughput.
@@
Figure 5 systematically quantify the performance degradation of chunked prefill of a 16k-token sequence prefill even without hybrid batching decode requests. For 1k chunk size, a progressive 10% drop in compute efficiency (from 71% to 61%) across successive chunks is witnessed in Figure 5(a), which falls substantially below the 77% achievable peak. This under-utilization stems from redundant KV cache reloading in chunked attention, which also causes the final chunk’s processing time to 1.9 than that of the initial chunk. Consequently, per-chunk latency scales linearly by chunk counts and increases total prefill latency by 1.13 compared to unchunked execution. While a larger chunk size of 2k partially mitigates utilization drops from -18% to -7%, the average per-chunk latency increases by 1.86, significantly diminishing the TPOT improvements that motivated chunked prefill. This fundamental tension between maintaining high hardware utilization and minimizing TPOT creates an intractable optimization in chunked prefill.
2.3.3. Existing Optimizations